### v1

Abstract
Timely and accurate lymphoma diagnosis is essential for guiding cancer treatment. Standard diagnostic practice combines hematoxylin and eosin (HE)-stained whole slide images with immunohistochemistry, flow cytometry, and molecular genetic tests to determine lymphoma subtypes, a process requiring costly equipment, skilled personnel, and causing treatment delays. Deep learning methods could assist pathologists by extracting diagnostic information from routinely available HE-stained slides, yet comprehensive benchmarks for lymphoma subtyping on multicenter data are lacking.
In this work, we present the first multicenter lymphoma benchmarking dataset covering four common lymphoma subtypes and healthy control tissue. We systematically evaluate five publicly available pathology foundation models (H-optimus-1, H0-mini, Virchow2, UNI2, Titan) combined with attention-based (AB-MIL) and transformer-based (TransMIL) multiple instance learning aggregators across three magnifications ($10\times$, $20\times$, $40\times$). On in-distribution test sets, models achieve multiclass balanced accuracies exceeding $80\%$ across all magnifications, with all foundation models performing similarly and both aggregation methods showing comparable results. The magnification study reveals that $40\times$ resolution is sufficient, with no performance gains from higher resolutions or cross-magnification aggregation. However, on out-of-distribution test sets, performance drops substantially to around $60\%$, highlighting significant generalization challenges. To advance the field, larger multicenter studies covering additional rare lymphoma subtypes are needed. We provide an automated benchmarking pipeline to facilitate such future research.
keywords:
Multicenter Lymphomas Benchmark, Multiple Instance Learning, Whole Slide Images, Pathology Foundation Models.
1 Introduction
Cancer is one of the deadliest diseases and remains an insurmountable obstacle to advance the quality and expectancy of life all over the world [bray2021ever]. Lymphoma is a type of blood cancer that originates in the lymphatic system, which is a critical part of human body’s immune system. It specifically arises from lymphocytes, white blood cells that play a key role in defending the body against infections. Lymphomas are broadly classified into two main categories [lewis2020lymphoma]: Hodgkin lymphoma (HL) and non-Hodgkin lymphoma (NHL), with each category having numerous subtypes. The diagnosis of lymphoma involves a combination of clinical evaluation, medical imaging, and most importantly, biopsy of the affected tissue. The biopsy is examined under a microscope (i.e., digitized as gigapixel HE (Hematoxylin and Eosin) stained whole slide images), and additional tests like immunohistochemical (IHC) stains, flow cytometry, cytogenetic, and molecular analysis help to determine the specific subtype of lymphoma [lewis2020lymphoma]. These auxiliary tests require costly equipment, expensive reagents, and trained personnel. Treatment varies depending on the lymphoma subtype, stage, and other factors such as the patient’s overall health. Common treatment options [lewis2020lymphoma] include chemotherapy, radiation therapy, targeted therapy, immunotherapy, and stem cell transplantation.
Histopathology plays a central role in clinical medicine for tissue-based diagnostics and in biomedical research as a basis for understanding cellular and molecular mechanisms of disease by microscopic evaluation of morphological changes [campanella2023computational]. Advances in histology slide scanning enable the digitization of entire histology slides into a whole slide image (WSI), paving the way for computer-aided diagnosis. In contrast to IHC, flow cytometry, and molecular genetic tests, HE-stained WSIs are inexpensive and widely available. Recent studies [vorontsov2024foundation-virchow1, zimmermann2024virchow2, chen2024towards, hoptimus1, ding2025titan, h0-mini] have shown rapid progress in pathology foundation models (FMs) on diverse large unlabeled WSI datasets. These models learn transferable representations that can be adapted to a wide range of downstream pathology tasks with relatively less annotated data. Among the most prominent pathology FMs are: Virchow2 [zimmermann2024virchow2], UNI2 [chen2024towards], H-optimus-1 [hoptimus1], Titan [ding2025titan], and H0-mini [h0-mini]. For downstream pathology tasks, Multiple Instance Learning (MIL) has been applied as a key methodology in computational pathology [ilse2018attention, campanella2019clinical, sadafi2020attention, shao2021transmil], where labels are assigned at the “bag” level (WSI), while learning is performed from individual instances (patches) within each bag. MIL is particularly well-suited for histopathology, where detailed annotations are often difficult to obtain.
Figure 1: A: Our multicenter lymphoma benchmark datasets comprises 5 classes : Follicular Lymphoma (FL), Chronic Lymphocytic Leukemia (CLL), Mantle Cell Lymphoma (MCL), Diffuse Large B-Cell Lymphoma (DLBCL) and healthy controls (NEG), with 999999 HE WSIs from 609609 individuals. B: Thumbnails of exemplary CLL WSIs. C: Random patches of $256\times 256$ pixels at $40\times$ magnification.
Earlier studies [vrabac2021dlbcl, 2010autocls, vorontsov2024foundation-virchow1] have demonstrated strong performance in distinguishing a limited set of lymphoma types, such as Diffuse Large B-Cell Lymphoma (DLBCL) and Follicular Lymphoma (FL) from HE-stained WSI. Their clinical utility remains limited due to narrow subtype coverage. In this work, we present first multicenter lymphoma benchmark covering four common subtypes simultaneously in Europe and healthy control tissue (Figure 1). Moreover, we provide an automated benchmarking MIL pipeline for external evaluation (Figure 2). Our main contributions are the following:
  • We present a first multicenter lymphoma benchmark for the four most frequent subtypes and healthy controls.
  • We provide a deep multiple instance learning pipeline with five publicly available pathology foundation models for validation studies.
  • We evaluate attention-based MIL models on in-distribution and out-of-distribution lymphoma datasets at three different resolutions.
Figure 2: A deep multiple instance learning pipeline from whole slide images acquisition to interpretability. A: Multiple whole slide images (WSIs) are obtained per patient after tissue sectioning, HE staining, and digital scanning. Thousand of the patches are extracted from each WSI by segmenting the tissue regions. B: The extracted patches are passed to a pretrained Encoder model (frozen) to extract feature vectors. The stacked feature vectors are fed to attention based aggregation layer (trainable) to aggregate patch-level information into slide-level representations, which are used to make the final lymphoma subtyping prediction. In addition, for each subtype, we extract attention scores to visualize important regions of the WSI.
2 Methodology
The deep learning pipeline for our experiments consists of lymphoma whole slide imaging, tissue segmentation and patch extraction, feature extraction, and lymphoma subtype classification (Figure 2). In the following, we describe each component in detail.
2.1 Lymphoma subtypes
Figure 1 shows the data used in our experiments. The lymphoma subtypes in the datasets are Chronic Lymphocytic Leukemia (CLL), Follicular Lymphoma (FL), Mantle Cell Lymphoma (MCL), and Diffuse Large B-Cell Lymphoma (DLBCL). In addition, the dataset contains WSIs showing only healthy tissue sections, which are used as a control class and designated as negative (NEG). CLL is a B-cell neoplasm that originates in the bone marrow and is characterized by the progressive accumulation of mature-appearing lymphocytes, which expand into the peripheral blood and infiltrate lymphoid organs, ultimately impairing normal organ function [ghia2007chronic]. FL arises from germinal-center B-cells—centrocytes and centroblasts that proliferate aberrantly and form characteristic follicular or follicle-like structures [xerri2016heterogeneity]. MCL develops from mantle-zone B-cells and typically presents with generalized lymphadenopathy or splenomegaly; it is defined by the expansion of cyclin-D1–driven malignant lymphocytes and displays an aggressive clinical course [schieber2018current]. DLBCL represents a high-grade B-cell lymphoma in which antibody-producing lymphocytes undergo malignant transformation, leading to rapidly enlarging masses in lymph nodes or extranodal sites due to uncontrolled proliferation [morton2006lymphoma]. Figure 3 shows the morphology of the four lymphoma classes in the Munich dataset.
Figure 3: Five lymphoma classes morphology overview in Munich dataset. Five randomly selected patches from a heterogenous WSI of each class at a size of $256\times 256$ pixels with a magnification of $40\times$, at which individual cells become recognisable.
2.1.1 Munich
[…]
2.1.2 Kiel
[…]
2.1.3 Augsburg
[…]
2.1.4 Erlangen
[…]
2.2 Tissue segmentation and patch extraction
Tissue segmentation filters out background areas in WSIs, reducing the number of irrelevant patches and speeding up non-overlapping patch extraction. Traditional techniques typically rely on image-processing heuristics such as binary or Otsu thresholding, which require manual parameter tuning and often fail to generalize across stains. These approaches also struggle to distinguish true tissue from common artifacts, including pen marks and bubbles. Trident [zhang2025standardizing] is a recent alternative for robust and stain-agnostic tissue segmentation and patch extraction. After tissue segmentation, the patching step divides the tissue-containing regions into $N$ non-overlap image patches with a size of $256\times 256$ pixel at specific magnification for later processing by a feature encoder (Figure 2A). We can represent the resulting patches as $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{N}\right\},\mathbf{x}_{i}\in[0,255]^{256\times 256\times 3}$, where $\mathbf{X}$ is called a bag for a WSI, and $\mathbf{x}_{i}$ is called the i-th instance.
2.3 Feature extraction
Due to the high dimensionality of a WSI, $\mathbf{X}$ usually consists of a large number of instances and cannot be processed by the downstream aggregation models. Thus, we first pass patches to a pretrained encoder to yield patch features by compressing the high dimensional visual information into a one dimensional feature vector. Lets suppose $f_{\psi}$ is a pretrained encoder with parameters $\psi$, for each patch $\mathbf{x}_{i}\in\mathbf{X}$ the corresponding feature vector is obtained as $\mathbf{h}_{i}=f_{\psi}\left(\mathbf{x}_{i}\right)$. The WSI is represented by a bag of instance features $\mathbf{H}=\left\{\mathbf{h}_{1},\ldots,\mathbf{h}_{N}\right\}$, which is passed as an input to the aggregation layer (see Figure 2B). In our pipeline, the pretrained encoders can either be used to extract patch-level or slide-level features.
2.4 Lymphoma subtype classification
We use the deep multiple instance learning (MIL) framework for the trainable aggregation layer to classify the bag $\mathbf{H}$ of instance features into the five classes of our lymphoma dataset. The aggregation layer takes as input a bag $\mathbf{H}$ of $N$ instance features and outputs a patient latent vector $\mathbf{z}=\sum_{k=1}^{N}a_{k}\mathbf{h}_{k}$, where $a_{k}$ is the attention score. Finally, the patient latent vector $\mathbf{z}$ is passed to classifier to obtain predicted class scores $\mathbf{\hat{Y}}$. During training, the cross-entropy loss is computed between the predicted WSI class label $\mathbf{\hat{Y}}$ and ground-truth label $\mathbf{Y}\in\left\{1,\ldots,C\right\}$.
3 Experimental Results
3.1 Preprocessing
Our HE-stained WSIs contain pen marks and artifacts, i.e., blurring, compression, water bubbles. In our pipeline, we use Trident [zhang2025standardizing] for both tissue segmentation and patch extraction; it employs a DeepLabV3-based segmentation model that provides more robust and stain-agnostic tissue detection and segmentation, and further extract non-overlapping patches from a WSI.
3.2 Evaluation
We split the patient-wise for five-fold cross-validation with train, validation, and test sets with a ratio of $80\%$, $10\%$, and $10\%$, respectively, to assess the performance of trained MIL models. To prevent data leakage and overfitting, all images of one patient are assigned to the same fold. The training set is shuffled to introduce slight variability, while the validation and test sets remain unshuffled to ensure consistent and reliable evaluation.
The model’s performance is evaluated using the Area Under the ROC Curve (AUC), Macro F1-score, and Balanced Accuracy (BACC) metrics. These metrics are calculated on both the validation and test datasets to provide a comprehensive assessment of classification effectiveness. Additionally, confusion matrices are analyzed to interpret the distribution of prediction errors.
Table 1: Magnification study: Five-fold cross-validation results for lymphoma subtype classification on Munich dataset at different magnification level, $40\times$, $20\times$, and $10\times$. Performance is reported in terms of Area Under ROC (AUC), Macro F1-score, and balanced accuracy (BACC).
Magnification Encoder
Embedding
Dim
Aggregator AUC F1-score BACC
ResNet50
1024 AB-MIL 0.89 $\pm$0.03 0.66 $\pm$0.06 0.67 $\pm$0.06
TransMIL 0.85 $\pm$0.02 0.58 $\pm$0.02 0.59 $\pm$0.03
TransMIL + BEL 0.87 $\pm$0.02 0.62 $\pm$0.04 0.63 $\pm$0.03
UNI2
(Patch-level)
1536 AB-MIL 0.95 $\pm$0.01 0.79 $\pm$0.02 0.79 $\pm$0.03
$40\times$ TransMIL 0.94 $\pm$0.01 0.75 $\pm$0.04 0.75 $\pm$0.04
TransMIL + BEL 0.96 $\pm$0.01 0.80 $\pm$0.03 0.80 $\pm$0.03
Titan
(Slide-level)
768 Linear 0.95 $\pm$0.01 0.78 $\pm$0.02 0.78 $\pm$0.01
AB-MIL 0.96 $\pm$0.01 0.83 $\pm$0.06 0.82 $\pm$0.05
TransMIL 0.96 $\pm$0.01 0.82 $\pm$0.03 0.82 $\pm$0.03
TransMIL + BEL 0.96 $\pm$0.01 0.82 $\pm$0.04 0.81 $\pm$0.04
ResNet50
1024 AB-MIL 0.89 $\pm$0.02 0.63 $\pm$0.04 0.64 $\pm$0.04
TransMIL 0.81 $\pm$0.04 0.50 $\pm$0.04 0.51 $\pm$0.04
TransMIL + BEL 0.84 $\pm$0.02 0.58 $\pm$0.02 0.58 $\pm$0.03
UNI2
(Patch-level)
1536 AB-MIL 0.95 $\pm$0.02 0.77 $\pm$0.03 0.77 $\pm$0.03
$20\times$ TransMIL 0.94 $\pm$0.01 0.79 $\pm$0.04 0.79 $\pm$0.04
TransMIL + BEL 0.96 $\pm$0.01 0.83 $\pm$0.03 0.83 $\pm$0.03
Titan
(Slide-level)
768 Linear 0.95 $\pm$0.01 0.77 $\pm$0.02 0.77 $\pm$0.02
AB-MIL 0.96 $\pm$0.01 0.82 $\pm$0.06 0.81 $\pm$0.07
TransMIL 0.96 $\pm$0.01 0.82 $\pm$0.04 0.82 $\pm$0.04
TransMIL + BEL 0.96 $\pm$0.01 0.82 $\pm$0.03 0.82 $\pm$0.04
ResNet50
1024 AB-MIL 0.85 $\pm$0.02 0.60 $\pm$0.07 0.61 $\pm$0.06
TransMIL 0.83 $\pm$0.04 0.52 $\pm$0.06 0.54 $\pm$0.06
TransMIL + BEL 0.82 $\pm$0.04 0.51 $\pm$0.04 0.53 $\pm$0.04
UNI2
(Patch-level)
1536 AB-MIL 0.95 $\pm$0.01 0.80 $\pm$0.04 0.80 $\pm$0.04
$10\times$ TransMIL 0.95 $\pm$0.01 0.80 $\pm$0.04 0.80 $\pm$0.04
TransMIL + BEL 0.96 $\pm$0.02 0.81 $\pm$0.05 0.81 $\pm$0.05
Titan
(Slide-level)
768 Linear 0.94 $\pm$0.01 0.75 $\pm$0.04 0.75 $\pm$0.04
AB-MIL 0.95 $\pm$0.02 0.79 $\pm$0.04 0.79 $\pm$0.04
TransMIL 0.94 $\pm$0.01 0.77 $\pm$0.04 0.78 $\pm$0.04
TransMIL + BEL 0.95 $\pm$0.01 0.80 $\pm$0.04 0.80 $\pm$0.04
3.3 Implementation details
We use several pretrained vision encoders for feature extraction, ResNet50 (baseline) [he2016resnet50], H-optimus-1 [hoptimus1], H0-mini [h0-mini], Virchow2 [zimmermann2024virchow2], UNI2 [chen2024uni], Titan [ding2025titan]. Each model independently process the WSI patches, generating s-dimensional embeddings encoding the visual representation of each WSI.
The MIL models are trained under a weakly supervised learning setting, as only slide-level labels are available and no detailed annotations for individual patches. We use the MIL architecture, employing three variants: attention-based deep MIL (AB-MIL) [ilse2018attention], transformer-based MIL (TransMIL) [shao2021transmil], and TransMIL + BEL [sens2023bel]. These models aggregate features extracted from individual instances (patches) within each WSI to generate a single, comprehensive representation of the entire slide. By learning to assign different weights to instances, the architectures can capture the most relevant morphological features contributing to the classification task.
The MIL models are trained for 50 epochs. Each training step involves a forward pass to generate predictions, computation of the cross-entropy loss by comparing predictions with ground-truth slide-level labels, and parameter updates via backpropagation using the AdamW optimizer, which combines weight decay with adaptive learning rate adjustments. To support convergence, a learning rate schedule with a warm-up phase of 10 epochs is used, linearly increasing the learning rate from 0 to 0.0001. After the warm-up, a cosine decay scheduler gradually reduces the learning rate down to 1e-6. Simultaneously, the weight decay, initially set to 0.04, are gradually increased up to 0.4. A momentum factor of 0.9 is also applied to accelerate optimization.
Table 2: Comparison of publicly available feature encoders: Five-fold cross-validation results for lymphoma subtype classification on Munich and Kiel lymphoma datasets at 40x magnification. Performance is reported in terms of AUC, F1-score, and balanced accuracy (BACC).
IID Test
Dataset Encoder
Embedding
Dim
Aggregator AUC F1-score BACC
Munich ResNet50 1024 AB-MIL 0.89 $\pm$0.03 0.66 $\pm$0.06 0.67 $\pm$0.06
H-optimus-1 1536 0.96 $\pm$0.01 0.79 $\pm$0.03 0.79 $\pm$0.03
H0-mini 1536 0.95 $\pm$0.01 0.79 $\pm$0.02 0.78 $\pm$0.01
Virchow2 2560 0.96 $\pm$0.01 0.77 $\pm$0.04 0.77 $\pm$0.04
UNI2 1536 0.95 $\pm$0.01 0.79 $\pm$0.02 0.79 $\pm$0.03
Titan 768 0.96 $\pm$0.01 0.83 $\pm$0.06 0.82 $\pm$0.05
Kiel ResNet50 1024 AB-MIL 0.85 $\pm$0.05 0.64 $\pm$0.08 0.65 $\pm$0.04
H-optimus-1 1536 0.97 $\pm$0.01 0.81 $\pm$0.09 0.81 $\pm$0.09
H0-mini 1536 0.92 $\pm$0.05 0.66 $\pm$0.15 0.69 $\pm$0.13
Virchow2 2560 0.96 $\pm$0.02 0.72 $\pm$0.11 0.74 $\pm$0.09
UNI2 1536 0.96 $\pm$0.03 0.82 $\pm$0.07 0.82 $\pm$0.06
Titan 768 0.96 $\pm$0.02 0.83 $\pm$0.06 0.81 $\pm$0.05
Table 3: Out-of-distribution generalization: Five-fold cross-validation results for lymphoma subtype classification on external datasets at $40\times$ magnification. UNI2 and Titan are used as feature encoders and pretrained aggregators (AB-MIL and TransMIL) on Munich dataset. Performance is reported in terms of AUC, F1-score, and balanced accuracy (BACC).
OOD Test
Encoder
OOD
Dataset
Aggregator AUC F1-score BACC
UNI2 Kiel AB-MIL 0.88 $\pm$0.05 0.61 $\pm$0.06 0.62 $\pm$0.04
TransMIL 0.86 $\pm$0.06 0.58 $\pm$0.06 0.60 $\pm$0.04
Augsburg AB-MIL 0.58 $\pm$0.07 0.45 $\pm$0.05 0.43 $\pm$0.04
TransMIL 0.47 $\pm$0.05 0.41 $\pm$0.05 0.40 $\pm$0.04
Erlangen AB-MIL 0.75 $\pm$0.05 0.37 $\pm$0.07 0.59 $\pm$0.18
TransMIL 0.90 $\pm$0.05 0.75 $\pm$0.35 0.83 $\pm$0.24
Titan Kiel AB-MIL 0.88 $\pm$0.05 0.43 $\pm$0.08 0.48 $\pm$0.08
TransMIL 0.84 $\pm$0.04 0.45 $\pm$0.10 0.50 $\pm$0.08
Augsburg AB-MIL 0.52 $\pm$0.04 0.30 $\pm$0.03 0.44 $\pm$0.04
TransMIL 0.47 $\pm$0.05 0.29 $\pm$0.05 0.44 $\pm$0.04
Erlangen AB-MIL 0.78 $\pm$0.05 0.35 $\pm$0.09 0.53 $\pm$0.20
TransMIL 0.86 $\pm$0.05 0.60 $\pm$0.25 0.72 $\pm$0.18
Figure 4: Confusion matrix comparison on in-distribution Munich and out-of-distribution testsets: IID testset, Munich, and external OOD test: Kiel, Augsburg, and Erlangen. The MIL models perform strongest on Munich, and better results on Kiel and Erlangen, while lower perfomance on Augsburg due to noisy IHC stained labels.
3.4 Quantitative results
For lymphoma subtype prediction, we assessed model performance across varying magnifications, multiple patch-level and slide-level feature encoders, and both in-distribution and out-of-distribution (OOD) test sets.
Magnification study (Table 1): On the in-distribution Munich test set, we compared baseline ResNet50, the patch-level encoder UNI2, and the slide-level encoder Titan across 40×, 20×, and 10× magnifications. TransMIL combined with UNI2 or Titan is better than AB-MIL at 40× and 20×, with a slight performance decline observed at 10×. Despite this, AB-MIL remained the most computationally efficient option for training and inference compared to transformer-based MIL methods.
Comparison of publicly available feature encoders (Table 2): Using AB-MIL, we evaluated several open-source pathology feature encoders on the Munich and Kiel in-distribution cohorts. UNI2 and Titan achieved superior performance across both datasets, indicating their strong representation capacity for lymphoma subtyping.
Out-of-distribution generalization (Table 3): Models trained on the Munich dataset using UNI2 or Titan features were evaluated on three external cohorts: Kiel, Augsburg, and Erlangen. UNI2 paired with AB-MIL achieved the best overall performance on OOD Kiel and Augsburg in the multi-class setting. For the Erlangen dataset, which contains only CLL cases, UNI2 combined with TransMIL yielded the highest accuracy. Across all OOD evaluations, UNI2 + AB-MIL demonstrated the most robust and stable generalization behavior.
Confusion matrix comparison on in-distribution and out-of-distribution testsets: Figure 4 presents confusion matrices for UNI2 + AB-MIL on the in-distribution Munich test set and the three OOD cohorts. The model performed strongest on the Munich dataset, followed by solid results on Kiel and Erlangen. Performance on Augsburg was comparatively lower, particularly for CLL and FL, which we attribute to noisy labels—Augsburg subtype labels were derived from IHC staining without pathologist confirmation, leading to potential misclassifications.
Figure 5: Attention visualization on IID Munich testset. Highest attention regions correspond to morphology, while lower attention regions mostly correspond to normal tissue or artifacts patches.
3.5 Whole slide image attention visualization
The attention scores can be visualized as heatmaps to highlight diagnostically informative regions, areas assigned high attention, while de-emphasizing regions of low relevance, such as normal tissue or background artifacts. To interpret the spatial distribution of model attention across a WSI, we convert the attention scores corresponding to the predicted class into percentiles and map these normalized values back to their original coordinates on the slide. Fine-grained heatmaps are produced by extracting overlapping patches and averaging the attention values within the overlapping areas. In our workflow, we employ Trident [zhang2025standardizing] to generate WSI-level attention visualizations (Figure 5 and Figure 6) and to display the patches with the highest attention scores. We note that a high-attention score in CLL is given to non-lymphoma areas and in DLBCL to poorly preserved areas (Figure 5). This behavior might be avoidable when training the algorithm with more classes.
Figure 6: Attention visualization of CLL lymphoma subtype on OOD lymphoma testsets.
4 Limitations
Although this multicenter benchmark is comprehensive, several limitations remain. First, the substantial decrease in performance ($20\%$) on external test sets highlights unresolved out-of-distribution generalization challenges. Scanner-specific color variation, staining differences, and shifts in patient cohorts across clinical sites continue to affect model robustness, as reflected by the performance drop on OOD cohorts. Second, the Augsburg dataset contains subtype labels derived automatically from IHC staining without pathologist confirmation, introducing label noise that likely contributes to the lower classification accuracy observed in this cohort. Finally, some cohorts exhibit class imbalance, particularly for rare subtypes or limited-control classes, which may bias model training and lead to overconfident predictions for majority classes.
5 Conclusion
We present the first multicenter lymphoma benchmarking dataset comprising 999 HE-stained whole slide images across four common lymphoma subtypes (CLL, FL, MCL, DLBCL) and healthy control tissue from four German centers. Combined with an automated multiple instance learning pipeline, we systematically evaluated five state-of-the-art pathology foundation models and two aggregation architectures across multiple magnifications.
On in-distribution test sets, all five foundation models (H-optimus-1, H0-mini, Virchow2, UNI2, Titan) achieved comparable performance with balanced accuracies exceeding $80\%$, demonstrating that current pathology foundation models have reached a similar level of representation quality for lymphoma subtyping. Both attention-based (AB-MIL) and transformer-based (TransMIL) aggregators performed similarly, with AB-MIL being computationally more efficient. Our magnification study revealed that 40× resolution is sufficient for accurate classification, with no additional benefit from higher resolutions or cross-magnification aggregation strategies. Despite strong in-distribution performance, we observed substantial performance drops on out-of-distribution test sets (balanced accuracy around $60\%$), revealing significant generalization challenges. This degradation likely stems from scanner-specific color variations, site-specific staining protocols, and differences in patient populations across centers. The Augsburg cohort, where labels were derived automatically from IHC staining without pathologist confirmation, showed particularly poor performance, highlighting the critical importance of high-quality ground truth annotations for model development. Additionally, we did not assess computational requirements or inference times, which are crucial considerations for clinical deployment. In contrast to expensive and time-consuming IHC staining, flow cytometry, and molecular genetic tests, our approach demonstrates that inexpensive HE-stained WSIs can accurately determine common lymphoma subtypes when evaluated within the same clinical environment. An AI-augmented pathology workflow could guide initial screening to prioritize cases requiring expert review, potentially reducing turnaround times and allowing more efficient use of specialized diagnostic tests.
To advance lymphoma diagnostics toward clinical utility, larger multicenter studies with standardized protocols are essential. Future work should expand coverage to include rare lymphoma subtypes, incorporate stain normalization or domain adaptation techniques to improve cross-site generalization, and conduct prospective clinical validation studies. Integrating complementary modalities such as immunophenotyping data or clinical parameters may further enhance diagnostic accuracy. Ultimately, such systems could assist pathologists in maximizing diagnostic yield from HE-stained sections while minimizing the number of auxiliary tests required for accurate diagnosis.
[…]
C.M. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 866411 & 101113551 & 101213822) and support from the Hightech Agenda Bayern.
Ethics statement
All experiments are conducted in accordance with the Declaration of Helsinki. The Technical University Munich (TUM) Ethics Committee approved the retrospective analysis of lymphoma data (Approval No. 79/20 S-KH).
Code availability statement
Our paper codes will be publicly available at https://github.com/RaoUmer/LymphomaMIL.